OpenStudy (anonymous):

URGENT!! PLEASE HELP! WILL MEDAL AND FAN! x+y=7 2x+3y=19 is (2,5) a solution to the linear equation? is (3,4) a solution?

OpenStudy (anonymous):

no 2,5 is

OpenStudy (anonymous):

I placed an equal sign where I think you meant to have typed one. Equation 1: x + y = 7 Equation 2: 2x + 3y = 19 Solve equation 1 for y: x + y = 7 Equation 3: y = 7 - x Substitute (7 - x) for y in equation 2: 2x + 3y = 19 2x + 3(7 - x) = 19 2x + 21 - 3x = 19 21 - x = 19 -x = -2 x = 2 Substitute (2) for x is equation 3: y = 7 - x y = 7 - (2) y = 5 So the solution is x = 2 and y = 5.

OpenStudy (anonymous):

Thanks! :)